DIRC Operations Report
Wednesday 2008/01/16 to Wednesday 2008/01/23
Oncall: Jennifer Prendki (Joe Schwiening on backup)
Fast Monitoring: Joe Schwiening and Jennifer Prendki
Luminosities this week
PEP peak lumi: ~10.0 x 1033
(last week: ~9.9 x 1033)
PEP delivery: 3.4 fb-1
BaBar integrated lumi: 3.22 fb-1
Run 7 total delivery: 13.95 fb-1
Run 7 total BaBar recorded: 13.66 fb-1 total,
including 0 fb-1 off-peak data and including 13.04
fb-1 on the Y(3S)
PEP activities/issues this week
- The week-end was successful (shift and 24hr lumi records was broken twice).
- A power dipped during the owl shift Tuesday, and
PEP had hard time struggling to recover.
- HER bend magnets down by 10 MeV for better tuning around the 3S CM energy.
- One big issue this week was the E_CM ajustment to maximize the luminosity. E_CM has been ajusted many times by about 1 MeV because PEP keeps sliding off the top of the peak. This is seen in the HadB ratios.
BaBar activities/issues this week
- On Friday, a BABAR magnet quench was triggered by a power dip.
- The difficult background conditions causes a lot of hard
time for many subsystems.
- Today, there was a lot of background while the currents were low. SVT, DCH, DRC and EMC were affected. This was a period with high dead times and
- On Tuesday, a second power dip occured but BABAR recovered quickly.
DIRC activities/issues this week
- The DIRC was not affected by any of the power dips.
- Jose made some tests to understand the shoulders seen on the
timing distributions on fast mon plots.
- This problem started on Sunday 13, with run 78206
- run 78259 had the distribution very bad
- no trickle injection for both beams --> run 78313
- trickle injection for only the LER --> run 78315
- trickle injection for only the HER --> run 78319
- Conclusion : the problem was due to bad trickle injection in
the LER. The background was only showing up on trigger lines #1
and #8 which are pure DCT triggers. Chris O'Grady has made some
tests with the master crates to make sure that there is no problem
with dataflow. Matt Weaver has shown that the bad timing
distributions are coming from pulses with less than 50
The data quality doesn't seem to be affected after
reconstruction. Even run 78259 looks definitely fine.
- Jose un-inhibited SIAM0 channel0 (the channel monitoring) the
chiller on/off. Since the new flow sensor was installed on Jan9,
this channel has been very stable. The new sensor is working fine.
- Some Wiener crates minor alarms occured on Thursday evening
at 7:40 PM, but these were just glitches.
- I (Jennifer) was automatically paged at 8:00 AM on Friday
because of a HV trip in sector 2, channel 3 - Jose was paged for
the same problem the day previous day. Trip can be seen in run
#78437 and #78444. The threshold was moved from 2295 uA to
2345 uA because the splashes seen by the other channels are spikes
of about 60uA ... No new alarm since then ...
- Jose gave the DIRC monthly report at the 3:45 meeting this afternoon.
You can find his presentation here :
See DIRC logbook for details
Lumi and Background this week
- Joe and Jennifer rotating oncall and backup duties.